# Separating semigroup of genus 4 curves

### v1

@@
In the present paper we describe the separating semigroups of all genus 4 curves. For the proofs we consider the canonical embedding of $C$ into a quadric $X$ in $\mathbb{P}^{3}$ and apply Abel’s theorem to 1-forms obtained as Poincaré residues at $C$ of certain meromorphic 2-forms on $X$.
@@ 1. Introduction
A necessary and sufficient condition for $C$ to be separating is the existence of a separating morphism $f:C\to\mathbb{P}^{1}$, that is a morphism such that $f^{-1}(\mathbb{RP}^{1})=\mathbb{R}C$. The restriction of a separating morphism to $\mathbb{R}C$ is a covering over $\mathbb{RP}^{1}$. If we fix the numbering of connected components $c_{1},\dots,c_{r}$ of $\mathbb{R}C$, we may consider the sequence $d(f)=(d_{1},\dots,d_{r})$ where $d_{i}$ is the covering degree of $f$ restricted to $c_{i}$. Kummer and Shaw [2] defined the separating semigroup of $C$ as
@@
It is shown in [2] that $\text{\rm Sep}(C)=\mathbb{N}^{g+1}$ if $C$ is an $M$-curve of genus $g$ (i.e. $\mathbb{R}C$ has $g+1$ connected components) and $\text{\rm Sep}(C)$ is $\mathbb{N}^{2}$ (resp. $\mathbb{N}^{3}$ or $2+\mathbb{N}_{0}$) if $C$ is a separating curve of genus 1 (resp. 2). The separating semigroups of hyperelliptic curves (of any genus) and of curves of genus 3 are computed in [3]. A much simpler proof for curves of genus 3 is given in [4, Remark 3.3].
In the present paper we compute the separating semigroup of all genus 4 curves (this result was announced in [4, Remark 3.6]). Let $C$ be a separating curve of genus 4. If $C$ is an $M$-curve, then $\text{\rm Sep}(C)=\mathbb{N}^{5}$ (see above). If $C$ is hyperelliptic but not an $M$-curve, then $\text{\rm Sep}(C)=\{2\}\cup(4+\mathbb{N}_{0})$ (see [3]).
@@
When $X$ is an ellipsoid or a hyperboloid, all rigid isotopy classes of smooth real sextic curves $C$ of genus 4 on $X$ (i.e. the connected components of the space of such curves) are described in [1]. The same arguments can be easily adapted to the case when $X$ is a quadratic cone (see also the footnote in [5, Lemma 23]). Representatives of all the rigid isotopy classes of separating non-maximal curves up to automorphisms of $X$ are depicted in Figure 1. Four of them are realizable as a small perturbation of three plane sections. The other two are perturbations of a plane section and a section by a thin cylinder around a line. Arrows in Figure 1 represent complex orientations, i.e., the boundary orientations induced from one of the halves of $C\setminus\mathbb{R}C$.
@@
Let $C$ be a separating curve of genus $4$ on a real quadric $X$. We number the components according to Figure 1 (arbitrarily if $C$ is an $M$-curve). Then $\text{\rm Sep}(X)$ is as in Table 1.
@@
The rest of the paper is devoted to the proof of Theorem 1. It is based on the techniques proposed in [4, §3].
Till the end of the paper, $X$ and $C$ are as in Theorem 1 and $C$ is not an $M$-curve (Theorem 1 for $M$-curves is proven in [2, Thm. 1.7]). We denote the number of connected components of $\mathbb{R}C$ by $r$ and the number of ovals of $C$ by $l$. As in the definition of $\text{\rm Sep}(C)$, the connected components of $\mathbb{R}C$ are denoted by $c_{1},\dots,c_{r}$ (numbered according to Figure 1).
@@ 2. Main lemmas
Let $D=D_{0}+D_{1}$ be a real plane section of $X$ where $D_{0}$ is the component of $D$ of even multiplicity (in our case $D_{0}$ is non-empty only when $X$ is the quadratic cone and $D=D_{0}$ is a double generator). The divisor $C-D$ belongs to the canonical class of $X$, thus it is the divisor of some holomorphic 2-form $\Omega_{D}$ on $X$. It defines a “chess-board” orientation on $\mathbb{R}X\setminus(\mathbb{R}C\cup\mathbb{R}D_{1})$, i.e., an orientation which changes when crossing $\mathbb{R}C\cup\mathbb{R}D_{1}$. We define the $D$-orientation of $\mathbb{R}C\setminus\mathbb{R}D$ as the boundary orientation induced by the “chess-board” orientation of $\mathbb{R}X\setminus(\mathbb{R}C\cup\mathbb{R}D)$. Let $\omega_{D}$ be the Poincaré residue of $\Omega_{D}$. Then the $D$-orientation coincides with the orientation defined by $\omega_{D}$ in the sense that $\omega_{D}(v)>0$ for $v\in T(\mathbb{R}C)$ if and only if the $D$-orientation is positive on $v$. Similarly to the complex orientations, the $D$-orientation is defined up to simultaneous reversing on all components of $\mathbb{R}C$.
@@
Let $p_{1},\dots,p_{n}$ and $q_{1},\dots,q_{n}$ be pairwise distinct points on $C$. Suppose that the divisors $P=p_{1}+\dots+p_{n}$ and $Q=q_{1}+\dots+q_{n}$ are linearly equivalent. Then there exist smooth paths $p_{j}:[0,t_{0}]\to C$, $p_{j}(0)=p_{j}$, $t_{0}>0$, such that $P_{t}=\sum p_{j}(t)\in|P|$ for each $t\in[0,t_{0}]$, and the derivative $p^{\prime}_{j}(0)$ is nonzero for each $j=1,\dots,n$.
@@
Case 3. $p_{j}\in D\cap D^{\prime}$ and $P\subset D\cup D^{\prime}$ (see the two bottom pictures in Figure 2). Each of $D$, $D^{\prime}$ passes through two points of $P-p_{j}$, hence $D^{\prime\prime}$ crosses $D\cup D^{\prime}$ only at points of $P$ because $D\cdot D^{\prime\prime}=D^{\prime}\cdot D^{\prime\prime}=2$. Choosing $D^{\prime}$ sufficiently close to $D$, we may assume that both $D^{\prime}$ and $D^{\prime\prime}$ are $\Cal{C}^{1}$-close to $D$, in particular, $D^{\prime\prime}$ goes monotonically from the left to the right (according to Figure 2). Easy to see that this is impossible under condition that $D^{\prime\prime}$ passes through all points of $P-p_{j}$ and crosses $D\cup D^{\prime}$ only at the points of $P$.
@@
By applying the same arguments with $D^{\prime}$ instead of $D$ we conclude that in all cases with $r=3$, the complex orientation on the vectors $v_{j}$ have the same sign whenever the corresponding points $p_{j}$ belong to the same connected component of $\mathbb{R}C$. In the case $r=1$ (the rightmost case in Figure 2), these arguments give only that the complex orientation has the same sign on the vectors $v_{2},v_{3},v_{4}$ (here we number the points $p_{j}$ from the left to the right). However, by applying the same arguments to the plane sections through $p_{2},p_{4},p_{5}$ (resp. $p_{1},p_{2},p_{3}$), we obtain that the complex orientation has the same sign on $v_{1}$ and $v_{3}$ (resp. on $v_{4}$ and $v_{5}$).
Therefore, for $0<t\ll t_{0}$, the divisors $D$ and $D_{t}$ are interlacing (see [2, §2.1]), i.e., each component of $\mathbb{R}C\setminus D$ contains exactly one point of $D_{t}$. Then (see [2, Prop. 2.11]) the meromorphic function on $C$ whose divisor is $D-D_{t}$ defines a separating morphism $C\to\mathbb{P}^{1}$. Hence we have $\{(1,3,1),(1,2,2)\}\subset\text{\rm Sep}(C)$ when $r=3$, and we have $5\in\text{\rm Sep}(C)$ when $r=1$. Since the divisor $P$ is non-special, the result follows from [2, Prop. 3.2]. ∎
@@
Let $X$ be a quadratic cone or hyperboloid and $L$ a straight line such that $\mathbb{R}L\subset\mathbb{R}X$. Then $L$ cannot have non-empty intersection with two ovals of $C$.
@@
Any curve on $X$ intersects an oval at an even number of points counting multiplicities. Thus the result follows from the fact that $L\cdot C=3$. ∎
@@ 3.1. Proof for ellipsoids
The pencil of planes through a point of $c_{1}$ and a point of $c_{3}$ defines a separating morphism which realizes $(1,2,1)\in\text{\rm Sep}(X)$.
@@ 3.2. Proof for quadratic cones
A pencil of plane sections passing through a real line $L$ defines a morphism $f:X\to\mathbb{P}^{1}$. If $L$ is a generator of $X$, then $f$ realizes $(1,1,1)\in\text{\rm Sep}(X)$. If $L$ meets $c_{1}$ and $c_{3}$ but not $c_{2}$, then $f$ realizes $(1,2,1)\in\text{\rm Sep}(X)$. ∎
@@
Let $L$ be a real line close the rotation axes of $X$ and not passing through the apex. Let $D$ be the section of $X$ by a real plane passing through $L$. Then (1) holds by Lemma 2.3, and we also have $(2,2,1)+\mathbb{N}_{0}^{3}\subset\text{\rm Sep}(C)$ by symmetry. ∎
@@
Suppose that $C$ is as in Figure 1, i.e. it is a perturbation of a plane section and a thin cylinder whose axes is linked with $c_{2}$. Then the pencil of plane sections passing through a line $L$ intersecting $c_{1}$ and $c_{3}$ realizes $(1,2,1)\in\text{\rm Sep}(X)$. As we pointed out in the introduction, any other curve with $(r,l)=(3,2)$ on $X$ is obtained from this model curve by a continuous deformation. Such a deformation can be followed by a simultaneous continuous deformation of the line $L$ intersecting $c_{1}$ and $c_{3}$. During the deformation, $L$ cannot become a generator of the cone by Lemma 2.4. Thus the line $L$ remains to be linked with $c_{2}$, hence the pencil of planes through $L$ always defines the same element of $\text{\rm Sep}(X)$. ∎
@@ 3.3. Proof for hyperboloids
$\pi_{A}:\mathbb{R}X\to\mathbb{R}A$ is a 3-fold covering (see (3)), hence it realizes $3\in\text{\rm Sep}(X)$.
@@
Thus $\Cal{D}$ realizes $4\in\text{\rm Sep}(X)$.
@@
Suppose that there exists a separating morphism $f:X\to\mathbb{P}^{1}$ realizing $(d_{1},1,d_{3})$. The projection $\pi_{A}:\mathbb{R}C\to\mathbb{R}A$ is not a covering. Hence there exists a fiber (we may assume that it is $A$) which has one real intersection point with $C$. Let $A\cap C=\{p\}$. Then $p\in c_{2}$, hence $A\cap(c_{1}\cup c_{2})=\varnothing$. Let $P=f^{-1}(f(p))$ and $D=A+B$. Then $D$ cannot pass through all points of $P$ by Lemma 2.4, and we obtain a contradiction with Lemma 2.1. ∎
@@
Let $D_{0}$ be the section of $X$ by a real plane passing through the line shown in Figure 1 and such that $[\mathbb{R}D_{0}]=[c_{2}]$ in $H_{1}(\mathbb{R}X)$. Then Lemma 2.3 implies (1) and, by symmetry, (2) for the model curve $C_{0}$. Let us choose continuous paths $\{p_{t}\}$ and $\{q_{t}\}$ so that $p_{t}\in c_{t,1}$, $q_{t}\in c_{t,3}$ for all $t$ and $p_{0},q_{0}\in\mathbb{R}D_{0}$. By Lemma 2.4 the line $p_{t}q_{t}$ is not contained in $X$ for all $t$. Hence by Lemma 3.7 there exists a continuous family of irreducible plane sections $\{D_{t}\}$ such that $p_{t},q_{t}\in\mathbb{R}D_{t}$ and $[\mathbb{R}D_{t}]=[c_{2}]$ for each $t$. Then the order of the points of $\mathbb{R}D_{t}\cap\mathbb{R}C_{t}$ does not change during the deformation, hence we may apply Lemma 2.3 to $C$. ∎

### v2

@@
In the present paper we describe the separating semigroups of all genus 4 curves. For the proofs we consider the canonical embedding of $C$ into a quadric $X$ in $\mathbb{P}^{3}$ and apply Abel’s theorem to 1-forms on $C$ obtained as Poincaré residues of certain meromorphic 2-forms.
support: A part of the work on the paper was done during author’s visit to Geneva supported by the Swiss National Science Foundation project 200400
@@ 1. Introduction
A necessary and sufficient condition for $C$ to be separating is the existence of a separating morphism $f:C\to\mathbb{P}^{1}$, that is a morphism such that $f^{-1}(\mathbb{RP}^{1})=\mathbb{R}C$. The restriction of a separating morphism to $\mathbb{R}C$ is a covering over $\mathbb{RP}^{1}$. If we fix a numbering of the connected components $c_{1},\dots,c_{r}$ of $\mathbb{R}C$, we may consider the sequence $d(f)=(d_{1},\dots,d_{r})$ where $d_{i}$ is the covering degree of $f$ restricted to $c_{i}$. Kummer and Shaw [2] defined the separating semigroup of $C$ as
@@
It is shown in [2] that $\text{\rm Sep}(C)=\mathbb{N}^{g+1}$ if $C$ is an $M$-curve of genus $g$ (i.e. $\mathbb{R}C$ has $g+1$ connected components) and $\text{\rm Sep}(C)$ is $\mathbb{N}^{2}$ (resp. $\mathbb{N}^{3}$ or $2+\mathbb{N}_{0}$) if $C$ is a separating curve of genus 1 (resp. 2). The separating semigroups of hyperelliptic curves of any genus and of curves of genus 3 are computed in [3]. A much simpler proof for curves of genus 3 is given in [4, Remark 3.3], and in §4 we also give a proof for hyperelliptic curve, which is essentially the same as in [3] but exposed from the point of view proposed in [4].
In the present paper we compute the separating semigroup of all genus 4 curves (this result was announced in [4, Remark 3.6]). Let $C$ be a separating curve of genus 4. If $C$ is an $M$-curve, then $\text{\rm Sep}(C)=\mathbb{N}^{5}$ (see above). If $C$ is hyperelliptic but not an $M$-curve, then $\text{\rm Sep}(C)=\{2\}\cup(4+\mathbb{N}_{0})$ (see [3] and §4).
@@
When $X$ is an ellipsoid or a hyperboloid, all rigid isotopy classes of smooth real sextic curves $C$ of genus 4 on $X$ (i.e. the connected components of the space of such curves) are described in [1]. The same arguments can be easily adapted to the case when $X$ is a quadratic cone (see also the footnote in [5, p. 14]). Representatives of all the rigid isotopy classes of separating non-maximal curves up to automorphisms of $X$ are depicted in Figure 1. Four of them are realizable as a small perturbation of three plane sections. The other two are perturbations of the union of a plane section and a section by a thin cylinder around a line. The arrows in Figure 1 represent complex orientations, i.e., the boundary orientations induced from one of the halves of $C\setminus\mathbb{R}C$.
@@
Let $C$ be a separating curve of genus $4$ on a real quadric $X$. We number the components according to Figure 1 (arbitrarily if $C$ is an $M$-curve). Then $\text{\rm Sep}(C)$ is as in Table 1.
@@
Sections 2 and 3 are devoted to the proof of Theorem 1. It is based on the techniques proposed in [4, §3]. In Section 4 we give a proof of [3, Theorem 2] similar to the proof of Theorem 1 of the present paper.
Till the end of Section 3, $X$ and $C$ are as in Theorem 1 and $C$ is not an $M$-curve (Theorem 1 for $M$-curves is proven in [2, Thm. 1.7]). We denote the number of connected components of $\mathbb{R}C$ by $r$ and the number of ovals of $C$ by $l$. As in the definition of $\text{\rm Sep}(C)$, the connected components of $\mathbb{R}C$ are denoted by $c_{1},\dots,c_{r}$ (numbered according to Figure 1).
@@ 2. Main lemmas
Let $D=D_{0}+D_{1}$ be a real plane section of $X$ where $D_{0}$ is the component of $D$ of even multiplicity (in our case $D_{0}$ is non-empty only when $X$ is the quadratic cone and $D=D_{0}$ is a double generator). The divisor $D-C$ belongs to the canonical class of $X$, thus it is the divisor of some meromorphic 2-form $\Omega_{D}$ on $X$. It defines a “chess-board” orientation on $\mathbb{R}X\setminus(\mathbb{R}C\cup\mathbb{R}D_{1})$, i.e., an orientation which changes when crossing $\mathbb{R}C\cup\mathbb{R}D_{1}$. We define the $D$-orientation of $\mathbb{R}C\setminus\mathbb{R}D$ as the boundary orientation induced by the “chess-board” orientation of $\mathbb{R}X\setminus(\mathbb{R}C\cup\mathbb{R}D_{1})$. Let $\omega_{D}$ be the Poincaré residue of $\Omega_{D}$. Then the $D$-orientation coincides with the orientation defined by $\omega_{D}$ in the sense that $\omega_{D}(v)>0$ for $v\in T(\mathbb{R}C)$ if and only if the $D$-orientation is positive on $v$. Similarly to the complex orientations, the $D$-orientation is defined up to simultaneous reversing on all components of $\mathbb{R}C$.
@@
Let $p_{1},\dots,p_{n}$ and $q_{1},\dots,q_{n}$ be pairwise distinct points on $C$. Suppose that the divisors $P=p_{1}+\dots+p_{n}$ and $Q=q_{1}+\dots+q_{n}$ are linearly equivalent. Then there exist smooth paths $p_{j}:[0,t_{0}]\to C$, $p_{j}(0)=p_{j}$, $t_{0}>0$, such that $P_{t}:=\sum p_{j}(t)\in|P|$ for each $t\in[0,t_{0}]$, and the derivative $p^{\prime}_{j}(0)$ is nonzero for each $j=1,\dots,n$.
@@
Case 3. $p_{j}\in D\cap D^{\prime}$ and $P\subset D\cup D^{\prime}$ (see the two bottom pictures in Figure 2). Each of $D$, $D^{\prime}$ passes through two points of $P-p_{j}$, hence $D^{\prime\prime}$ crosses $D\cup D^{\prime}$ only at points of $P$ because $D\cdot D^{\prime\prime}=D^{\prime}\cdot D^{\prime\prime}=2$. Choosing $D^{\prime}$ sufficiently close to $D$, we may ensure that both $D^{\prime}$ and $D^{\prime\prime}$ are $\Cal{C}^{1}$-close to $D$, in particular, $D^{\prime\prime}$ goes monotonically from the left to the right (according to Figure 2). Easy to see that this is impossible under condition that $D^{\prime\prime}$ passes through all points of $P-p_{j}$ and crosses $D\cup D^{\prime}$ only at the points of $P$.
@@
By applying the same arguments with $D^{\prime}$ instead of $D$ we conclude that in all cases with $r=3$, the complex orientation on the vectors $v_{j}$ have the same sign whenever the corresponding points $p_{j}$ belong to the same connected component of $\mathbb{R}C$. In the case $r=1$ (the rightmost case in Figure 2), these arguments give only that the complex orientation has the same sign on the vectors $v_{2},v_{3},v_{4}$ (here we number the points $p_{j}$ from the left to the right). However, by applying the same arguments to the plane section through $p_{2},p_{4},p_{5}$ (resp. $p_{1},p_{2},p_{3}$), we obtain that the complex orientation has the same sign on $v_{1}$ and $v_{3}$ (resp. on $v_{4}$ and $v_{5}$).
Therefore, for $0<t\ll t_{0}$, the divisors $P$ and $P_{t}$ are interlacing (see [2, §2.1]), i.e., each component of $\mathbb{R}C\setminus P$ contains exactly one point of $P_{t}$. Then (see [2, Prop. 2.11]) the meromorphic function on $C$ whose divisor is $P-P_{t}$ defines a separating morphism $C\to\mathbb{P}^{1}$. Hence we have $\{(1,3,1),(1,2,2)\}\subset\text{\rm Sep}(C)$ when $r=3$, and we have $5\in\text{\rm Sep}(C)$ when $r=1$. Since the divisor $P$ is non-special, the result follows from [2, Prop. 3.2]. ∎
@@
Suppose that $X$ is a quadratic cone or hyperboloid, $l=2$, and $L$ is a real line on $X$. Then $L$ cannot have non-empty intersection with both ovals of $C$.
@@
Any real curve on $X$ intersects an oval at an even number of points counting multiplicities. Thus the result follows from the fact that $L\cdot C=3$. ∎
@@ 3.1. Proof for ellipsoids
The pencil of planes through a point of $c_{1}$ and a point of $c_{3}$ defines a separating morphism which realizes $(1,2,1)\in\text{\rm Sep}(C)$.
@@ 3.2. Proof for quadratic cones
A pencil of plane sections passing through a real line $L$ defines a morphism $f:X\to\mathbb{P}^{1}$. If $L$ is a generator of $X$, then $f$ realizes $(1,1,1)\in\text{\rm Sep}(C)$. If $L$ meets $c_{1}$ and $c_{3}$ but not $c_{2}$, then $f$ realizes $(1,2,1)\in\text{\rm Sep}(C)$. ∎
@@
Let $L$ be a real line close to the rotation axis of $X$ and not passing through the apex. Let $D$ be the section of $X$ by a real plane passing through $L$. Then (1) holds by Lemma 2.3, and we also have $(2,2,1)+\mathbb{N}_{0}^{3}\subset\text{\rm Sep}(C)$ by symmetry. ∎
@@
Suppose that $C$ is as in Figure 1, i.e. it is a perturbation of a plane section and a thin cylinder whose axis is linked with $c_{2}$. Then the pencil of plane sections passing through a line $L$ intersecting $c_{1}$ and $c_{3}$ realizes $(1,2,1)\in\text{\rm Sep}(C)$. As we pointed out in the introduction, any other curve with $(r,l)=(3,2)$ on $X$ is obtained from this model curve by a continuous deformation. Such a deformation can be followed by a simultaneous continuous deformation of the line $L$ intersecting $c_{1}$ and $c_{3}$. During the deformation, $L$ cannot become a generator of the cone by Lemma 2.4. Thus the line $L$ remains to be linked with $c_{2}$, hence the pencil of planes through $L$ always defines the same element of $\text{\rm Sep}(C)$. ∎
@@ 3.3. Proof for hyperboloids
$\pi_{A}:\mathbb{R}X\to\mathbb{R}A$ is a 3-fold covering (see (3)), hence it realizes $3\in\text{\rm Sep}(C)$.
@@
Thus $\Cal{D}$ realizes $4\in\text{\rm Sep}(C)$.
@@
Suppose that there exists a separating morphism $f:X\to\mathbb{P}^{1}$ realizing $(d_{1},1,d_{3})$. The projection $\pi_{A}:\mathbb{R}C\to\mathbb{R}A$ is not a covering. Hence there exists a fiber (we may assume that it is $A$) which has one real intersection point with $C$. Let $A\cap C=\{p\}$. Then $p\in c_{2}$, hence $A\cap(c_{1}\cup c_{3})=\varnothing$. Let $P=f^{-1}(f(p))$ and $D=A+B$. Then $D$ cannot pass through all points of $P$ by Lemma 2.4, and we obtain a contradiction with Lemma 2.1. ∎
@@
Let $D_{0}$ be the section of $X$ by a real plane passing through the line shown in Figure 1 and such that $[\mathbb{R}D_{0}]=[c_{2}]$ in $H_{1}(\mathbb{R}X)$. Then Lemma 2.3 implies (1) and, by symmetry, (2) for the model curve $C_{0}$. Let us choose continuous paths $\{p_{t}\}$ and $\{q_{t}\}$ so that $p_{0},q_{0}\in\mathbb{R}D_{0}$ and $p_{t}\in c_{t,1}$, $q_{t}\in c_{t,3}$ for all $t$. By Lemma 2.4 the line $p_{t}q_{t}$ is not contained in $X$ for all $t$. Hence by Lemma 3.7 there exists a continuous family of irreducible real plane sections $\{D_{t}\}$ such that $p_{t},q_{t}\in\mathbb{R}D_{t}$ and $[\mathbb{R}D_{t}]=[c_{2}]$ for each $t$. Then the mutual arrangement of $\mathbb{R}D_{t}$ and $\mathbb{R}C_{t}$ does not change during the deformation, hence we may apply Lemma 2.3 to $C$. ∎
4. Hyperelliptic curves
The main results of [3] can be also proved using the approach of the present paper. One of them (a description of $\text{\rm Sep}(C)$ when genus$(C)=3$) is already reproved in this way in [4]. Here we give a new proof of the other.
[…]
([3, Thm. 1].) Let $C$ be a non-maximal hyperelliptic curve of genus $g\geq 1$. Set $m=\lfloor(g+1)/2\rfloor$. Then
$\text{\rm Sep}(C)=\begin{cases}\big{(}(1,1)\mathbb{N}\big{)}\cup\big{(}(m,m)+\mathbb{N}_{0}^{2}\big{)},&\text{if $g$ is odd,}\\ (2\mathbb{N})\cup(g+\mathbb{N}_{0}),&\text{if $g$ is even.}\end{cases}$ (5)5( 5 )
Demonstration Proof
The curve $C$ is defined by the equation $y^{2}=F(x)$ where $F$ is a real polynomial of degree $22g+$ 2 positive on $\mathbb{R}$. It can be embedded to a real Hirzebruch surface $X$ of degree $g+1$ (the fiberwise compactification of the line bundle $\Cal{O}(g+1)$) so that the hyperelliptic projection $\pi:C\to\mathbb{P}^{1}$ is the restriction of the fibration $X\to\mathbb{P}^{1}$. The restriction $\pi|_{\mathbb{R}C}$ is a two-fold covering over $\mathbb{RP}^{1}$. It is trivial if $g$ is odd and non-trivial if $g$ is even. We fix an affine chart $U\subset X$ with coordinates $(x,y)$ such that $C$ and $\pi$ are given by $y^{2}=F(x)$ and $(x,y)\mapsto x$ respectively. Then $\mathbb{R}C\cup U$ has two connected components $c_{1}$ and $c_{2}$. Each of them cuts transversally all fibers of $\pi$. We have $K_{X}+C\sim(g-1)F$ where $F$ is a fiber of $\pi$. Denote the semigroup in the right hand side of (5) by $S$.
Let us show that $\text{\rm Sep}(C)\subset S$. Let $f:C\to\mathbb{P}^{1}$ be a separating morphism and $P$ is its fiber contained in $U$. Let $P_{i}=P\cap c_{i}$ and $n_{i}=\#P_{i}$, $i=1,2$. Without loss of generality we may assume that $n_{1}\leq n_{2}$. Recall that $m=\lfloor(g+1)/2\rfloor$. Suppose that $n_{1}<m$. Let $D_{0}$ be the union of $m-1$ fibers of $\pi$ such that $P_{1}\subset D_{0}$ and $D_{0}\setminus\pi^{-1}(P_{1})$ is disjoint from $P$. If $g$ is odd, we set $D=2D_{0}$; if $g$ is even, we set $D=2D_{0}+F$, where $F$ is the fiber at infinity, i.e., the fiber of $\pi$ not contained in $U$. In both cases we have $D-C\sim K_{X}$ and we may consider the corresponding meromorphic 2-form $\Omega_{D}$ on $X$ and its Poincaré residue $\omega_{D}$ on $C$, which is a holomorphic 1-form; $\omega_{D}$ defines a $D$-orientation on $\mathbb{R}C\setminus\text{\rm supp}D$ (cf. §2). Then $\omega_{D}$ vanishes at the points of $P_{1}$ and the $D$-orientation coincides with the complex orientation on $c_{2}\setminus\text{\rm supp}D$. Hence $P_{2}\subset D_{0}$ by [4, Thm. 3.2], and hence $P_{2}\subset\pi^{-1}(P_{1})$. By symmetry we also have $P_{1}\subset\pi^{-1}(P_{2})$ and hence $n_{1}=n_{2}$, which implies $\text{\rm Sep}(C)\subset S$.
[…]
Now let us prove the inverse inclusion $S\subset\text{\rm Sep}(C)$. The hyperelliptic projection realizes $2\in\text{\rm Sep}(C)$ or $(1,1)\in\text{\rm Sep}(C)$ (depending on the parity of $g$), hence the semigroup generated by this element is contained in $\text{\rm Sep}(C)$. Thus, by [2, Prop. 3.2], it is enough to realize $g\in\text{\rm Sep}(C)$ or $(m,m)\in\text{\rm Sep}(C)$ (depending on the parity of $g$) by a separating morphism with non-special fibers. Let $p_{i}=(x_{i},y_{i})$, $i=1,\dots,g$, be points on $C$ such that $x_{1}<x_{2}<\dots<x_{g}$ and sign$(y_{i})=(-1)^{i}$, i.e., $p_{1},p_{3},\dots$ are on $c_{1}$ and $p_{2},p_{4},\dots$ are on $c_{2}$. Then the divisor $P=p_{1}+\dots+p_{g}$ is non-special on $C$. Let us show that $P$ is realizable as a fiber of a separating morphism. We proceed as in the proof of Lemma 2.3. By Riemann–Roch Theorem, $\dim|P|=1$. Let $P_{t}=p_{1}(t)+\dots+p_{g}(t)$, $P_{0}=P$, be a deformation of $P$ in $|P|$. The linear system $|P|$ does not have base points. Indeed, if $p_{i}(t)$ is constant, then $P^{*}_{t}:=P_{t}-p_{i}(t)\sim P^{*}:=P-p_{i}$, hence $P^{*}_{t}+\tau(P^{*})\sim P^{*}+\tau(P^{*})$, where $\tau$ is the hyperelliptic involution, but this contradicts the fact that $P^{*}+\tau(P^{*})\in|K_{C}|$ and that each element of $|K_{C}|$ is invariant under $\tau$. Let $D_{i}$, $i=1,\dots,g-1$, be the union of $g-2$ fibers of $\pi$ passing through all points of $P$ except $p_{i}$ and $p_{i+1}$. Then $D_{i}\in|K_{X}+C|$ and we consider the $D_{i}$-orientation on $C\setminus\text{\rm supp}(D_{i})$. It coincides with the complex orientation at one of the points $p_{i}$, $p_{i+1}$ and it is opposite at the other point (see Figure 3). Hence (cf. the end of the proof of Lemma 2.3), the complex orientation has the same sign on the tangent vectors $p^{\prime}_{i}(0)$ and $p^{\prime}_{i+1}(0)$. This is true for all $i=1,\dots,g-1$, hence the divisors $P$ and $P_{t}$, $0<t\ll 1$, are interlacing and the result follows from [2, Prop. 3.2].